Q. Choose the correct answer.
Q. Which of the following is the sum of infinite series ∑_(n=0)^∞▒〖(-1)〗^n 3^n r^n?
a. 1/(1+3r) for -1/3<r<1/3
b. . 1/(1-3r) for -1/3<r<1/3
. 3/(1-3r) for -1/3<r<1/3
d. ∞
1
Expert's answer
2019-07-05T10:41:33-0400
"1/(1-x)=1+x+x^2+...+x^n=\sum_{n=0} ^\infin x^n, |x|<1"

- the geometric series.

Then


"1\/(1+x)=1\/(1-(-x))="

"=1-x+x^2+...+(-1)^nx^n=\sum_{n=0} ^\infin (-1)^nx^n, |x|<1"

If

"x=3r"

then


"\sum_{n=0}^\infin (-1)^n 3^nr^n=1/(1+3r)"

for

"|3r|<1 \Rightarrow - 1/3<r<1/3."